# Restricted Path Characteristic Function Determines the Law of Stochastic Processes

### v2

@@ 1.4. The generalised moment problem; radius of convergence
  • Signature — $X=S(\gamma)$ for some path $\gamma$, e.g., of bounded $p$-variation for $21\leq p<$ 2;
@@
By virtue of Proposition 1.3, in the above cases Question 1.4 specialises to the following
Question 1.5.
@@
Notice that certain conditions must be imposed on $X$ to get an affirmative answer to Question 1.4. Even in the simplest case of real sequences (Example 1.2), to solve for each $x_{n}$ via (7), the radius of convergence of the power series in (6) must be non-zero.
[…]
@@
The radius of convergence of a tensor $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, written as $\rho(X)$, is the radius of convergence of the (complex) power series
@@
${\mathscr{E}}_{R}:=\left\{X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}:\,\rho(X)\geq R\right\},\qquad\widehat{{\mathscr{E}}}:=\bigcup_{R>0}{\mathscr{E}}_{R},\qquad\text{ and }{\mathscr{E}}_{\infty}:=\bigcap_{R>0}{\mathscr{E}}_{R}.$
The radius of convergence $\rho(X)$ of a tensor $X$ is closely related to its generating function $\Phi_{X}$ as in Definition 1.1, as shown by the following
@@
    If $X\in{\mathscr{E}}_{\infty}$ (i.e., $\rho(X)=\infty$), then the generating function $\Phi_{X}$ in (1.1) is well defined on the domain ${\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)$.
@@
    For each $R>0$, if $X\in{\mathscr{E}}_{R}$ (i.e., $\rho(X)\geq R$), then $\Phi_{X}$ is well defined on the subdomain $\left\{M\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right):\,\text{the operator norm of $M$}<R\right\}$.
@@
    $X\in{\mathscr{E}}_{\infty}$ if and only if it lies in the closure of ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$ with respect to the coarsest topology for which the following holds: for any normed algebra $\mathcal{A}$ and any algebra homomorphism $M\in{\rm Hom}\left(\mathbb{R}^{d},\mathcal{A}\right)$, the canonical extension ${\widetilde{M}}:{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}\to\mathcal{A}$ is continuous.
@@
Clearly $(2)\Rightarrow(1)$. To see (2), we simply note that $\|\widetilde{M}(\pi_{n}(X))\|\leq\|\pi_{n}(X)\|\cdot\|M\|_{\rm op}^{n}$ for each $n\in\mathbb{N}$, where $\|M\|_{\rm op}$ is the operator norm of $M$. Finally, (3) follows from [7, Corollary 2.5], with $\Psi$ chosen to be the singleton of the $\ell_{1}$-norm on $\mathbb{R}^{d}$ therein. ∎
@@ 1.5. Main Theorem
Our main result of this paper is to give an affirmative answer to Question 1.4, under the assumption $\rho(X)>0$ — equivalently, $X\in\widehat{{\mathscr{E}}}$. In fact, we shall obtain an explicit algorithm to solve for $X$ from the generating function $\Phi$:
@@
Given a mapping
$\Phi:\bigoplus_{k=0}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\to\bigoplus_{k=0}^{\infty}{\mathfrak{gl}(k,\mathbb{C})}$
graded by $k\in\mathbb{N}$. Suppose that $\Phi$ is the generating function of some $X\in\widehat{{\mathscr{E}}}$, i.e., a tensor with a nonzero radius of convergence. One can determine $X$ explicitly by the following procedure:
@@
${{\mathcal{M}}[W]}_{1},\ldots,{{\mathcal{M}}[W]}_{d}\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{so}}(k,\mathbb{R})\right)$
@@
Here, $p^{W}(i|t)$ is the location where $i$ appears for the $t^{\text{th}}$ time in $W=(w_{1},\ldots,w_{n})$, and $\sum^{\prime}_{t}$ means summation over $t$ from $1$ up to the total number of times that $i$ appears in $W$. The symbol $E^{a}_{b}$ designates the $k\times k$ matrix whose $(a,b)$-entry is $1$ and all the other entries are 00.
@@
An immediate consequence of Theorem 1.8 is the characteristic property of the generating function. That is, generating functions (viewed as the map $X\mapsto\Phi_{X}$) separate points over $\widehat{{\mathscr{E}}}$.
@@
For any $X\in\widehat{{\mathscr{E}}}$ with $\pi_{n}(X)\neq 0$ and any $k\geq n+1$, we can find $M\in{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$ such that $\Phi_{X}(M)\neq 0$. Here $\Phi_{X}$ is the generating function of $X$ as in Definition 1.1, and $\mathfrak{t}(k,\mathbb{R})$ is the space of tridiagonal antisymmetric $k\times k$ matrices as in Equation 1.
This result should be compared with Chevyrev–Lyons [7, Theorem 4.8], which ascertains that for any $X\in{\mathscr{E}}^{\infty}$ with $\pi_{n}(X)\neq 0$, we can find $M\in{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{sp}(m)\right)$ for some $m\geq\max\left\{2,\frac{n}{3}\right\}$ such that $\Phi_{X}(M)\neq 0$. Here $\mathfrak{sp}(m)$ is the Lie algebra of (the real form of the) symplectic group of $2m\times 2m$ real matrices, which contains $\mathfrak{t}(2m,\mathbb{R})$ as a subset. The proof in [7] is based on the polynomial identities in the numerical linear algebra literature [8, 9].
Given the above discussions on radius of convergence, we can say more about Question 1.5.
  • As is shown in Chevyrev–Lyons [7, Corollary 2.5 and §5], the signature of a $p$-rough path for any $p\geq 1$ has an infinite radius of convergence. Hence, its moment generating function uniquely determines the signature. In particular, for $p\in[1,2)$ the signature can be defined as in (2) in the sense of Young’s integrals, and the signature of any geometric rough path in $\mathbb{R}^{d}$ lies in $\mathcal{G}\left(\mathbb{R}^{d}\right)$, the set of group-like elements in tensor algebra. In this case, the uniqueness of unitary development holds.
  • The case for the expected signature is more complicated. The infinity of radius of convergence for the expected signature has been verified for several popular stochastic processes on a fixed time horizon, e.g., fractional Brownian motions with Hurst parameter in $[\frac{1}{2},1)$ ([21]), as well as certain Gaussian and Markovian rough paths ([7, Examples 6.7 and 6.8]). It is, nonetheless, considerably challenging to study processes up to a random time. The only known results in this regard, to the best of our knowledge, pertain to the Brownian motion up to the first exit time from a bounded domain $\Omega\subset\mathbb{R}^{d}$ (a.k.a. the “stopped Brownian motion”). Using PDE approaches pioneered by Lyons–Ni [18], it has been established in [1, 13] that the expected signature of stopped Brownian motions on $C^{2,\alpha}$-domains in $\mathbb{R}^{d}$ for $2\leq d\leq 8$ has a finite radius of convergence.
@@ 1.6. Applications to numerical analysis and machine learning
The theoretical justification of the PCF distance is essentially based upon the characteristic property of the (expected) unitary development. In Theorem 1.8 and its proof, we specify a reduced set of sparse linear maps in $\mathcal{L}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)$, under which the corresponding unitary development uniquely determines the unparameterised path. It motivates us to consider the “restricted PCF distance”, which characterises the law of time series while reducing the computation complexity compared with the PCF distance in [15]. This will be introduced in §6, with demonstration of its efficacy in hypothesis testing on the fractional Brownian motion.
@@ 1.7. Organisation
In §2, we introduce some notations that shall be used throughout this paper. Then, in §3, a combinatorial identity on derivatives of the unitary development of special matrices is presented. Our main Theorem 1.8 is proved in §4. The following section §5 is devoted to the special case of the (expected) signature. Next, in §6, we investigate the application of our main Theorem 1.8 to hypothesis testing, via the concept of restricted path characteristic function distance introduced therein. Some concluding remarks are given in §7. Appendix A presents the details of the numerical experiments on hypothesis testing.
@@ 2.1. Words
${\bf Word}(n):=\Big{\{}W:\,\text{$W$ is a word of length $n$ fromthe alphabet $\{1,\ldots,d\}$}\Big{\}}.$
@@ 2.2. Multiplicity and position functions
$\qquad\text{and }\quad p^{W}(i|t)=\emptyset\quad\text{ for all the other $i\in\{1,2,3,4\}$ and $t\in\left\{1,\ldots,r^{W}(i)\right\}$}.$
@@ 2.3. Matrices
$E^{a}_{b}\in{\mathfrak{gl}(k,\mathbb{C})}\quad\text{such that the $(a^{\prime},b^{\prime})$-entry is $\delta_{a}^{a^{\prime}}\delta^{b}_{b^{\prime}}$ for any $a^{\prime},b^{\prime}\in{\{1,\ldots,k\}}$}.$
That is, $E^{a}_{b}$ is the matrix whose $(a,b)$-entry is $1$ and all the other entries are 00.
@@
[…]
@@ 4. Determination of tensor from the development
Notice that the comments ensuing (18) explains the necessity of the assumption $X\in\widehat{{\mathscr{E}}}$. Also, by virtue of (20), nonzero contributions to the summation on the right-hand side of (22) arises only from the pairs $(I,{\overline{W}})$ such that $I\sim W$.
@@
The (1,n+1)-entry of $\text{The $(1,n+1)$-entry of }\left\langle\left\langle I,{\overline{W}}\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}=\begin{cases}1\qquad\text{if }I=W={\overline{W}},\\ 0\qquad\text{if otherwise},\end{cases}$
[…]
@@
Recall here that $E^{a}_{b}\in{\mathfrak{gl}(k,\mathbb{C})}$ are the elementary matrices. The multiplicity function $r^{W}(i)$ and position function $p^{W}(i|t)$ are as in Definition 2.1. Note too that ${{\mathcal{M}}[W]}_{i}(e_{j})$ have only nonzero entries (which are $\pm$) on the super- and sub-diagonal, i.e., at $(q,q\pm 1)$-coordinates.
@@
$\left\langle\left\langle I,{\overline{W}}\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}={{\mathcal{M}}[W]}_{i_{1}}\left(e_{{\overline{w}}_{1}}\right)\bullet\cdots\bullet{{\mathcal{M}}[W]}_{i_{n}}\left(e_{{\overline{w}}_{n}}\right)\qquad\text{where }I\sim W.$
@@
[…]
@@
$\lambda E^{1}_{2}\bullet E^{2}_{3}\bullet\cdots\bullet E_{n}^{n-1}\qquad\text{for some $\lambda\in\mathbb{R}$.}$ (29)
@@
For each {1,,n}, there is t{1,,rW(i)} such that $\text{For each $\ell\in{\{1,\ldots,n\}}$, there is $t\in\left\{1,\ldots,r^{W}(i_{\ell})\right\}$ such that $p^{W}\left(i_{\ell}|t\right)=\ell$}.$
@@
For each {1,,n}, the index i must occur at location  in the word $\text{For each $\ell\in{\{1,\ldots,n\}}$, the index $i_{\ell}$ must occur at location $\ell$ in the word $W$}.$
@@
A careful examination reveals that the term $-E^{p^{W}(i|t)+1}_{p^{W}(i|t)}$ in our expression (25) for ${{\mathcal{M}}[W]}_{i}$ is somewhat immaterial to the proof of Theorem 1.8. We include it only to ensure that ${{\mathcal{M}}[W]}_{i}$ takes value in the space ${\mathfrak{so}}(k,\mathbb{R})$ of anti-symmetric matrices. This leaves open the possibility for adapting our proof to Cartan developments into more general Lie groups.
@@
5. Determination of laws of stochastic processes
We now discuss the implications of our Main Theorem 1.8 on the problem of determining measures of stochastic processes from the expected development. Indeed, Theorem 1.8 identifies a new family of sparse matrix-valued linear mappings distinguishing the measures associated to path signatures, thus improving the classical result [7, Theorem 4.8] by Chevyrev–Lyons.
5.1. Separation of points
Recall from (1) the space of tridiagonal anti-symmetric matrices:
$\mathfrak{t}(k,\mathbb{R})$ :=$:=\bigg{\{}M\in\mathbb{R}^{k\times k}:\,\text{$M+M^{\top}=0$, andthe $(i,j)$-entry of $M$}$
$\qquad\qquad\qquad\qquad\qquad\text{is possibly nonzero only when$|i-j|=1$}\bigg{\}}.$
Also recall Corollary 1.9 (reproduced below):
Let $X\in\widehat{{\mathscr{E}}}$ be such that $\pi_{n}(X)\neq 0$ for some $n\in\mathbb{N}$. Then, for any $k\geq n+1$ one can find $M\in\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$ such that $\Phi_{X}(M)\neq 0$.
A direct consequence is as follows (the development map $\mathscr{D}_{\gamma}$ is defined in (5)):

### v3

@@ 1.4. Radius of convergence of a tensor
  • Signature — $X=S(\gamma)$ for some path $\gamma$, e.g., of bounded $p$-variation for $1\leq p<2$;
@@
By virtue of Proposition 1.4, in the above cases Question 1.5 specialises to the following:
Question 1.6.
@@
Notice that certain conditions must be imposed on $X$ to get an affirmative answer to Question 1.5. Even in the simplest case of real sequences (Example 1.3), to solve for each $x_{n}$ via (12), the radius of convergence of the power series in (11) must be non-zero.
The notion of radius of convergence can be extended to the elements of ${\mathcal{T}}\left(\mathbb{R}^{d}\right)$}. This shall play a central role in the investigation of Questions 1.5 &\& 1.6. Throughout this paper, to fix the ideas, we equip $\mathbb{R}^{d}$ with the $\ell_{1}$-norm, namely that $\|a\|:=\sum_{i=1}^{d}|a_{i}|$ for $a=(a_{1},\ldots,a_{d})^{\top}$, and equip $\left(\mathbb{R}^{d}\right)^{\otimes n}$ with the projective tensor norm for each $n=1,2,3,\ldots$, denoted by $\|\cdot\|$ again [^2] See [10, pp.136–139] and [22, p.3] for discussions on these norms. However, one may also define $\mathscr{E}$$_{\infty}$ as the completion of $\bigoplus_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n}$ under scalar dilations of any system of sub-multiplicative norms, as per [7, Remark 5.3].. Also recall the level-$n$ projection operator $\pi_{n}$ from (17). Then, we introduce:
@@
The radius of convergence (ROC) of a tensor $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, written as $\rho(X)$, is defined by the radius of convergence of the (complex) power series
@@
${\mathscr{E}}_{R}:=\left\{X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}:\,\sum_{n=0}^{\infty}\left\|\pi_{n}(X)\right\|R^{n}\text{ is convergent}\right\},\quad\widehat{{\mathscr{E}}}:=\bigcup_{R>0}{\mathscr{E}}_{R},\quad{\mathscr{E}}_{\infty}:=\bigcap_{R>0}{\mathscr{E}}_{R}.$
Note that if $X\in{\mathscr{E}}_{R}$, then $\rho(X)\geq R$. Also, ${\mathscr{E}}_{R}$ is a Polish space under the metric $\sum_{n=0}^{\infty}\left\|\pi_{n}(X)-\pi_{n}(Y)\right\|R^{n}$.
The radius of convergence $\rho(X)$ of a tensor $X$ is closely related to its generating function $\Phi_{X}$ defined in Definition 1.2, as shown by the following proposition.
@@
    If $X\in{\mathscr{E}}_{\infty}$, i.e., $\rho(X)=\infty$, then the generating function $\Phi_{X}$ in (13) is well defined on the domain ${\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)$.
@@
    If $X\in{\mathscr{E}}_{R}$, then $\Phi_{X}$ is well defined on $\left\{M\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right):\,\text{the operator norm of $M$}<R\right\}$.
@@
    $X\in{\mathscr{E}}_{\infty}$ if and only if it lies in the closure of $\bigoplus_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n}$ with respect to the coarsest topology for which the following holds: for any normed algebra $\mathcal{A}$ and any algebra homomorphism $M\in{\rm Hom}\left(\mathbb{R}^{d},\mathcal{A}\right)$, the canonical extension ${\widetilde{M}}:{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}\to\mathcal{A}$ is continuous.
@@
Clearly $(2)\Rightarrow(1)$. To see (2), we simply note that $\left\|\widetilde{M}(\pi_{n}(X))\right\|\leq\|\pi_{n}(X)\|\cdot\|M\|_{\rm op}^{n}$ for each $n\in\mathbb{N}$, where $\|M\|_{\rm op}$ is the operator norm of $M$. Finally, (3) follows from [7, Corollary 2.5], with $\Psi$ chosen to be the singleton of the $\ell_{1}$-norm on $\mathbb{R}^{d}$ therein. ∎
@@ 1.5. Main Theorem
Our main result of this paper provides an affirmative answer to Question 1.5, under the assumption $\rho(X)>0$ (equivalently, $X\in\widehat{{\mathscr{E}}}$). In fact, we establish an explicit algorithm to solve for $X$ from its generating function $\Phi_{X}$:
@@
Given the generating function (graded by $k\geq 1$)
$\Phi_{X}:\bigoplus_{k=1}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\longrightarrow\bigoplus_{k=1}^{\infty}{\mathfrak{gl}(k,\mathbb{C})}$
of some $X\in\widehat{{\mathscr{E}}}$, i.e., a tensor with a non-zero radius of convergence. Then, one can recover $X$ explicitly by the following procedure:
@@
${{\mathcal{M}}[W]}_{1},\ldots,{{\mathcal{M}}[W]}_{d}\in{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$
@@
Here, $p^{W}(i|t)$ is the location where $i$ appears for the $t^{\text{th}}$ time in $W=(w_{1},\ldots,w_{n})$, and $\sum^{\prime}_{t}$ means summation over $t$ from $1$ up to the total number of times that $i$ appears in $W$. The symbol $E^{a}_{b}$ designates the $k\times k$ matrix whose $(a,b)$-entry is $1$ and all the other entries are 0.
@@
An immediate consequence of Theorem 1.9 is the characteristic property of the generating function. That is, generating functions (viewed as the map $X\mapsto\Phi_{X}$) separate points in $\widehat{{\mathscr{E}}}$.
@@
For any $X\in\widehat{{\mathscr{E}}}$ with $\pi_{n}(X)\neq 0$ and any $k\geq n+1$, one can find $M\in{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)$ such that $\Phi_{X}(M)\neq 0$. Here $\Phi_{X}$ is the generating function of $X$ as in Definition 1.2, and $\mathfrak{t}(k,\mathbb{R})$ is the space of tridiagonal skew-symmetric $k\times k$ real matrices as in (3).
This result should be compared with Chevyrev–Lyons [7, Theorem 4.8], which ascertains that for any $X\in{\mathscr{E}}^{\infty}$ with $\pi_{n}(X)\neq 0$, one can find $M\in{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{sp}(m)\right)$ for some $m\geq\max\left\{2,\frac{n}{3}\right\}$ such that $\Phi_{X}(M)\neq 0$. Here $\mathfrak{sp}(m)$ is the Lie algebra of (the real form of the) symplectic group of $2m\times 2m$ real matrices, which contains $\mathfrak{t}(2m,\mathbb{R})$ as a subset. The proof in [7], based on polynomial identities in the numerical linear algebra literature [8, 9], is non-constructive by nature.
Given the above discussions on radius of convergence, we can say more about Question 1.6.
  • As shown in Chevyrev–Lyons [7, Corollary 2.5 and §5], the signature of a $p$-rough path for any $p\geq 1$ has an infinite radius of convergence. Hence, its moment generating function uniquely determines the signature. In particular, for $p\in[1,2[$ the signature can be defined as in (5) in the sense of Young’s integrals, and the signature of any geometric rough path in $\mathbb{R}^{d}$ lies in $\mathcal{G}\left(\mathbb{R}^{d}\right)$, the set of group-like elements in ${\mathcal{T}}\left(\mathbb{R}^{d}\right)$}. In this case, the uniqueness of unitary development holds.
  • The case for the expected signature is more complicated. The infinity of radius of convergence for the expected signature has been verified for several popular stochastic processes on a fixed time horizon, e.g., fractional Brownian motions with Hurst parameter in $\left[\frac{1}{2},1\right[$ ([23]), as well as certain Gaussian and Markovian rough paths ([7, Examples 6.7 and 6.8]). It is, nonetheless, considerably challenging to study processes up to a random time. The only known results in this regard, to the best of our knowledge, pertain to the Brownian motion up to the first exit time from a bounded domain $\Omega\subset\mathbb{R}^{d}$ (a.k.a. the “stopped Brownian motion”). Using PDE approaches pioneered by Lyons–Ni [20], it has been established in [1, 14] that the expected signature of stopped Brownian motions on $C^{2,\alpha}$-domains in $\mathbb{R}^{d}$ for $2\leq d\leq 8$ has a finite radius of convergence.
@@ 1.6. Applications to numerical analysis and machine learning
The theoretical justification of the PCF distance is essentially based upon the characteristic property of the (expected) unitary development. In Theorem 1.9 and its proof, we specify a reduced set of sparse linear maps in $\mathcal{L}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)$, under which the corresponding unitary development uniquely determines the unparameterised path. It motivates us to consider the “restricted PCF distance”, which characterises the law of time series while reducing the computation complexity compared with the PCF distance in [16]. This will be introduced in §6, with demonstration of its efficacy in hypothesis testing on fractional Brownian motion.
@@ 1.7. Organisation
In §2, we introduce some notations that shall be used throughout this paper. Then, in §3, a combinatorial identity on derivatives of the unitary development of special matrices is presented. Our Main Theorem 1.9 is proved in §4. The following section §5 is devoted to the theoretical implications of Main Theorem 1.9, i.e., the determination of measures on subspaces of ${\mathcal{T}}\left(\mathbb{R}^{d}\right)$} and path spaces. Finally, in §6, we investigate the application of Main Theorem 1.9 to hypothesis testing, via the concept of restricted path characteristic function distance introduced therein. Appendix A presents the details of the numerical experiments on hypothesis testing.
@@ 2.1. Words
[…]
@@ 2.2. Multiplicity and position functions
[…]
@@ 2.3. Matrices
[…]
That is, $E^{a}_{b}$ is the matrix whose $(a,b)$-entry is $1$ and all the other entries are 0.
@@
$\left\{E^{a}_{b}-E^{b}_{a}:\,1\leq a<b\leq k\right\}$ is a natural basis for ${\mathfrak{so}}(k,\mathbb{R})$.
@@ 4. Determination of tensor from the development
Notice that the comments ensuing (25) explains the necessity of the assumption $X\in\widehat{{\mathscr{E}}}$. Also, by virtue of (27), non-zero contributions to the summation on the right-hand side of (29) arises only from the pairs $(I,{\overline{W}})$ such that $I\sim W$.
@@
$\text{The $(1,n+1)$-entry of }\left\langle\left\langle I,{\overline{W}}\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}=\begin{cases}1\qquad\text{if }I=W={\overline{W}},\\ 0\qquad\text{if otherwise},\end{cases}$
 $\qquad\qquad\qquad\text{ for any $I,{\overline{W}}\in{\bf Word}(n)$ such that $I\sim W$}.$ (30)
@@
Recall here that $E^{a}_{b}\in{\mathfrak{gl}(k,\mathbb{C})}$ are the elementary matrices. The multiplicity function $r^{W}(i)$ and position function $p^{W}(i|t)$ are as in Definition 2.1. Note too that ${{\mathcal{M}}[W]}_{i}(e_{j})$ have only non-zero entries (which are $\pm$) on the super- and sub-diagonal, i.e., at $(q,q\pm 1)$-coordinates.
@@
[…]
@@
$\left\langle\left\langle I,I\right\rangle\right\rangle_{{{\mathcal{M}}[W]}}$ = $=\text{ linear combination of matrices of the form }E^{a_{1}}_{b_{1}}\bullet\cdots\bullet E^{a_{n}}_{b_{n}},$ (35)
@@
[…]
@@
[…]
@@
[…]
@@
A careful examination reveals that the term $E^{p^{W}(i|t)+1}_{p^{W}(i|t)}$ in our expression (32) for ${{\mathcal{M}}[W]}_{i}$ is somewhat immaterial to the proof of Theorem 1.9. We include it only to ensure that ${{\mathcal{M}}[W]}_{i}$ takes value in the space ${\mathfrak{so}}(k,\mathbb{R})$ of skew-symmetric matrices. This leaves open the possibility for adapting our proof to Cartan developments into more general Lie groups.
@@
5. Determination of measures over tensors of non-zero ROC
We now discuss implications of Theorem 1.9 to the determination of Radon measures on subspaces of the extended tensor algebra ${\mathcal{T}}\left(\mathbb{R}^{d}\right)$} and on the space of finite-$p$-variation paths.
5.1. Determination of measures on ${\mathscr{E}}_{R}$
Recall from (3) the space of tridiagonal skew-symmetric matrices $\mathfrak{t}(k,\mathbb{R})$. Also recall the definitions of the ROC of a tensor $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$ and of ${\mathscr{E}}_{R}$ (see Definition 1.7). The following version of the Stone–Weierstrass Theorem can be found, e.g., in [29, Exercise 7.14.79]:
Lemma 5.1.
Let $\mu$ and $\nu$ be Radon measures on a topological space $E$, and let $\mathcal{F}$ be a family of bounded continuous functions that separates points over $E$, contains the constant function $1$, and is closed under pointwise multiplication. If $\int_{E}f\,{\rm d}\mu=\int_{E}f\,{\rm d}\nu$ for every $f\in\mathcal{F}$, then $\mu=\nu$.
Fix any $R>0$. As in [7], we equip ${\mathscr{E}}_{R}$ with any Hausdorff topology such that each $\widetilde{M}\in{\mathcal{A}_{\mathfrak{t},R}}$ is continuous. We have shown in Theorem 1.9 that
[…]
separates points over ${\mathscr{E}}_{R}$. That is, for any $X_{1}\neq X_{2}$ in ${\mathscr{E}}_{R}$, there exists some $\widetilde{M_{0}}\in{\mathcal{A}_{\mathfrak{t},R}}$ such that $\widetilde{M_{0}}(X_{1})\neq\widetilde{M_{0}}(X_{2})$. Note that $\mathfrak{t}(k;\mathbb{R})$ is not an algebra: this motivates us to consider
${\bf Clos}\left[{\mathcal{A}_{\mathfrak{t},R}}\right]:=\text{ the smallest algebra containing }{\mathcal{A}_{\mathfrak{t},R}}.$
By an abuse of notation, we identify matrix algebras with the algebras of continuous functions of the matrix coefficients (denoted as $\mathcal{C}$ in [7, p.4058, §4]).
The result below follows immediately from Lemma 5.1. It together with Proposition 1.8 (3) directly implies [7, Corollary 4.12], which states that the Borel probability measures on the subspace of group-like elements $\mathcal{G}\left(\mathbb{R}^{d}\right)\subset{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$ are determined by $\mathcal{A}$ defined in (2).
Proposition 5.2.
Let $R>0$ and $\mathcal{B}\subseteq{\mathscr{E}}_{R}$ be a Borel set such that every $f\in{\bf Clos}\left[{\mathcal{A}_{\mathfrak{t},R}}\right]$ is bounded on $\mathcal{B}$. Suppose $\mu$ and $\nu$ are Radon measures on ${\mathscr{E}}_{R}$. If $\int_{\mathcal{B}}f\,{\rm d}\mu=\int_{\mathcal{B}}f\,{\rm d}\nu$ for every $f\in{\bf Clos}\left[{\mathcal{A}_{\mathfrak{t},R}}\right]$, then $\mu=\nu$ on $\mathcal{B}$.
Remark 5.3.
  1. (1)
    In Proposition 5.2, any compact subset of ${\mathscr{E}}_{R}$ is a valid choice of $\mathcal{B}$.
  2. (2)
    For the case $R=\infty$, $\mathscr{E}$$_{\infty}$ can be topologised as a Polish space (see [7, Corollary 2.4]). Hence, every finite Borel measure on $\mathscr{E}$$_{\infty}$ is Radon.
  3. (3)
    ${\mathscr{E}}_{R}$ is a Polish space under the metric $d_{R}(X,Y):=\sum_{n=0}^{\infty}\|\pi_{n}(X)-\pi_{n}(Y)\|R^{n}$ for any $0<R<\infty$.
  4. (4)
    The space of group-like elements $\mathcal{G}\left(\mathbb{R}^{d}\right)$ is a closed subspace of $\mathscr{E}$$_{\infty}$. The arguments in [7, §4] lead to the characterisation: ${\bf Clos}\left[{\mathcal{A}_{\mathfrak{t},R}}\right]\Big{|}_{\mathcal{G}\left(\mathbb{R}^{d}\right)}={\bf Clos}\left[{\mathcal{A}_{\mathfrak{t},R}}\big{|}_{\mathcal{G}\left(\mathbb{R}^{d}\right)}\right]$.
5.2. Determination of measures on path spaces
Fix $1\leq p<2$. Denote by $\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)$ the space of bounded-$p$-variation paths in $\mathbb{R}^{d}$ over $[0,T]$. Also denote its time-augmentation by
$\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right):=\left\{(\gamma_{t},t)_{t\in[0,T]}:\gamma\in\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d}\right)\right\}.$
Clearly $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)\subset\mathcal{V}_{p}\left([0,T],\mathbb{R}^{d+1}\right)$. Also, the signature maps $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$ injectively into $\mathcal{G}\left(\mathbb{R}^{d}\right)$, which is a closed subspace of $\mathscr{E}$$_{\infty}$ (see Remark 5.3).
Recall the definition of path development $\gamma\mapsto\mathscr{D}_{\gamma}$ from (9). As a direct consequence of Theorem 1.9, the path development separates points over $\tilde{\mathcal{V}}_{p}\left([0,T],\mathbb{R}^{d}\right)$.